Solve the equation √3sin𝜃−cos𝜃=0 for −180°≤𝑥≤180°


1
Expert's answer
2022-03-15T09:27:45-0400

"\sqrt{3} \sin\theta=\cos \theta;\\
\cfrac{\sin\theta}{\cos\theta}=\cfrac{1}{\sqrt3};\\
\tan\theta=\cfrac{1}{\sqrt3};\\
\theta=\arctan\cfrac{1}{\sqrt3}+180\degree\cdot n=30\degree+180\degree\cdot n, \ \ n\in Z;\\"

for "-180\degree\leq\theta\leq180\degree\ \ n=0" or "n=-1;"

"\theta_1=-150\degree,\ \ \theta_2=30\degree."